Meeting Notes for 30 March

Present: Joel, Ken, Liz, Sudha, Dave H (background)

Cosmic GPS

Lots of progress on fixing errant GPS coordinates on Cosmic detectors. Sudha reports that most errant DAQs are now covered. She, Mark and Ken are cleaning up the last few, including Tulsa, Elisa (Gatz?), (Matea?) Valley, 6665 in the middle of the Caribbean, and Wisconsin River Falls. Two in particular came up for discussion:
DAQ 6303

Originally issued to someone at Florida International University. Now appears to be in Mexico City (as of '09).

Ken's theory is that people at (Eberro?) mistook the number on the GPS unit as their DAQ number. Dave H seconds the theory.

This raises a potential problem: If true, correcting it will require reassigning the group's DAQ number. This means they might lose access to previously-uploaded data under the old DAQ number. Reassigning the old data to the new number might require "surgery," in Ken's words.

This isn't necessarily a problem at the moment, since the theory may be wrong and since the group might still have their data saved locally (they could then re-upload it under the new DAQ #). Nevertheless, it might be worthwhile to develop a general method for dealing with problems like this if they crop up in the future.

Liz brings up a few good questions to keep in mind when formulating a method, including
  • Has the data been used to make a plot? If so, we need to reassign without breaking the plot.
  • Is the data self-documenting, in the sense that DAQ number is included with all data files? If not, can it be and should it be in the future?
Ken reports that Tom told him the e-Labs strip lines from the raw data when loaded, but he doesn't remember about DAQ information. The full original data file is saved only locally, though.

Consensus is that it's best to get Mark's input on the larger DAQ#-changing problem, while Ken and Sudha keep investigating the situation with this particular detector.
DAQ 6139

This is showing up in the north Atlantic Ocean. Dave's sheet says it was returned to FNAL. Its last upload was 6/12/2008. We should look at the last data upload file to determine where it was at the time of upload. Mark and Sudha are working on it.


We had a couple of problems with the Drupal site this week (unresolved as of this annotation) that are probably related to shifting servers again a couple of weeks ago. Senem is working on them.

Liz hasn't had much response from Xeno Media about upgrading the site, and she doesn't think Marge has, either.

Summer CMS Workshops

Looking ahead to the summer, Liz asks about CMS Workshops. Ken has just started organizing them.
  • Shane has created a Google Doc for workshop scheduling. It can be shared with the development group to assist with the IT components of the workshops
  • The CMS e-Lab is ready, but Ken and Tom need to get the new data interface connected.
-- Main.JoelG - 2016-03-30


Topic revision: r1 - 2016-03-30, JoelG
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback